OpenStudy (anonymous):

36to40? A. 3 over 12 C. 20% D. 9 to 10

OpenStudy (sasogeek):

what is the question asking us to do?

OpenStudy (anonymous):

36 to 40 is the term and A, C and D are the definitions and i have to find the right one.

OpenStudy (callisto):

D

OpenStudy (sasogeek):

Callisto.. why?

OpenStudy (callisto):

36:40 = 36/40 = 9/10 = 9:10